The top and bottom margins of a poster are each 6 cm and the side margins are each 4cm. If the area of printed material on the poster is fixed at 384cm^3, find the dimensions of the poster with the smallest area?
Thankyou for your time.

Answer by stanbon(75887)

The top and bottom margins of a poster are each 6 cm and the side margins are each 4cm. If the area of printed material on the poster is fixed at 384cm^2, find the dimensions of the poster with the smallest area?
---------------------------
Assume the printed area is a square.
It's dimensions are each sqrt(384)
-------------------------------------
If you include the margins the dimensions are as follows:
length = sqrt(384)+12cm
width = sqrt(384) + 8cm
-------------
The total area of the poster is LW = 384 + 96 + 20sqrt(384)
= 480 + 20sqrt(384)
-----------------
The area of the margin is total area - printed material area
= 480 + 20sqrt(384) - 384 = 96+20sqrt(384) = 387.92 cm^2
==================
Cheers,
Stan H.